sir please explain hcv volume 1 chapter 14 Q10,Q11,Q12,Q14,Q22,Q23,Q24,Q26,Q27

Grade:12

1 Answers

Nirmal Singh.
10 years ago
Please send question one by one
Q 23 - Surface tension of water S = 0.075 J/m^2
radius r = 6 cm = 0.06 m
Area A = pie*r^2 = 3.14*(0.06^2) = 113.04 x10^-4 m^2
Syrface Energy U =S.A
U = 0.075*113.04*10^-4 = 8.478x10^-4 Joule
Thanks & Regards,
Nirmal Singh
Askiitians Faculty